# Experience Retrieval-Augmentation with Electronic Health Records Enables Accurate Discharge QA

### v1

@@ Abstract
To improve the reliability of Large Language Models (LLMs) in clinical applications, retrieval-augmented generation (RAG) is extensively applied to provide factual medical knowledge. However, beyond general medical knowledge from open-ended datasets, clinical case-based knowledge is also critical for effective medical reasoning, as it provides context grounded in real-world patient experiences. Motivated by this, we propose Experience Retrieval-Augmentation (ExpRAG) framework based on Electronic Health Record (EHR), aiming to offer the relevant context from other patients’ discharge reports. ExpRAG performs retrieval through a coarse-to-fine process, utilizing an EHR-based report ranker to efficiently identify similar patients, followed by an experience retriever to extract task-relevant content for enhanced medical reasoning. To evaluate ExpRAG, we introduce DischargeQA, a clinical QA dataset with 1,280 discharge-related questions across diagnosis, medication, and instruction tasks. Each problem is generated using EHR data to ensure realistic and challenging scenarios. Experimental results demonstrate that ExpRAG consistently outperforms a text-based ranker, achieving an average relative improvement of 5.2%, highlighting the importance of case-based knowledge for medical reasoning.
@@ 1 Introduction
Figure 1: An illustrative example of utilizing experience from relevant clinical cases to support medical decision.
@@
In light of this, we argue that, in addition to general factual concepts, clinical case-based knowledge is also crucial for effective medical reasoning. The intuition is that an experienced clinician often relies on past cases with similar conditions to guide diagnosis, treatment decisions, and discharge planning. In this study, we propose Experience Retrieval-Augmentation (ExpRAG) framework, leveraging a large-scale EHR database MIMIC-IV Johnson et al. (2023b) as its knowledge basis. Specifically, ExpRAG breaks down the retrieval process into two steps: (1) report ranking applies an EHR-based similarity measurement to identify patients with similar medical conditions, and (2) experience retrieval extracts problem-relevant content from these patients’ discharge reports, which serves as the case-based contextual knowledge for LLMs. The introduced EHR modality enables large-scale clinical experience retrieval, grounding the model’s reasoning in real-world clinical practices.
To evaluate ExpRAG’s capability in medical reasoning, we introduce DischargeQA  [^1] Uploading to https://physionet.org/ , please check updates on Github: https://github.com/jou2024/EXPRAG, a clinical dataset including 1,280 QA pairs dedicated to discharge-related problems. The dataset primarily includes three types of problems, simulating the discharge process from initial diagnosis to medication prescription and post-discharge instructions. For each problem, we follow the data structure of the discharge report and select the content preceding the question as the problem background to avoid label leakage. Additionally, we index the option candidates using EHR to generate contextually relevant options, ensuring a non-trivial and clinically meaningful challenge for the model.
We evaluate the performance of four different LLMs and compare ExpRAG with the text-based report ranker using DischargeQA. The results demonstrate the effectiveness of using EHR to retrieve relevant clinical experience, as it consistently improves the performance of LLM backbones and outperforms the text-based ranker with an average relative improvement of 5.2%. Overall, our main contributions are listed below:
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) RAG has become a key paradigm for overcoming the static knowledge limitations of LLMs by retrieving external information Gu et al. (2018); Petroni et al. (2019). Traditional RAG frameworks typically use dense retrieval methods to augment generative tasks Devlin et al. (2019); Xiong et al. (2021). While effective in general QA, these methods often lack domain specificity, which is critical in healthcare Lu et al. (2024). Recent advancements like ClinicalRAG Lu et al. (2024) and MIRAGE Xiong et al. (2024) address this by integrating structured EHR data and clinical notes for diagnosis and treatment planning. However, existing benchmarks primarily focus on isolated information retrieval Johnson et al. (2023a); Kweon et al. (2024), overlooking the complexities of reasoning over patient histories and similar cases. To bridge this gap, our work extends RAG by combining structured EHR data with discharge summaries, enabling experience-driven reasoning for more realistic and reliable medical QA.
Medical QA Benchmark EHRSQL (Lee et al., 2022) and DrugEHRQA (Bardhan et al., 2022) target structured data queries, with the former addressing SQL-based operations and the latter focusing on drug-related questions. EHRNoteQA (Kweon et al., 2024) and RadQA (Soni et al., 2022) leverage clinician-verified QA pairs from discharge summaries and radiology reports, while MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019) evaluate LLMs with questions from medical exams or PubMed articles. Discharge-summary-focused datasets like emrQA (Pampari et al., 2018) and CliniQG4QA (Yue et al., 2021) use discharge notes for QA tasks, and specialized datasets like RxWhyQA (Fan, 2019) and drug-reasoning QA (Moon et al., 2023) focus on specific question types like medication reasoning. Unlike these benchmarks, DischargeQA introduces an evaluation framework centered around the discharge process, simulating the clinical workflow from diagnosis inference to medication prescription and discharge instruction generation. Additionally, we leverage EHR to generate non-trivial, contextually relevant candidate options, providing a more challenging and realistic setup.
@@ 3 ExpRAG Framework
ExpRAG provides a comprehensive framework for retrieving relevant knowledge from the cohort, as shown in Figure 2. We first formulate the problem that ExpRAG aims to tackle and elaborate the two-step retrieval framework including an EHR-based report ranker.
@@ 3.1 Task Formulation
3.2 Coarse-to-Fine Retrieval Framework
@@
Report Ranking Before addressing a specific medical query, an intuitive assumption is that only patients with similar clinical histories, e.g., similar diseases or medications, can potentially provide meaningful guidance. In light of this, ExpRAG first employs a report ranker to efficiently discard unrelated cases and narrow down the candidate pool using the patient information:
@@
where $\mathcal{D}^{\prime}=\{D_{i}\}_{N^{\prime}\ll N}$ is a small subset of the selected discharge summaries. The ranker module will be scalable and enable effective utilization of patient context. Specifically, we introduce EHR as a knowledge basis to facilitate the patient-level similarity measurement, as presented in Section 3.3.
Experience Retrieval Based on the selected candidate pool, a sophisticated text retriever is capable of providing more accurate and dedicated clinical experience searching:
@@
Built on top of clinically relevant reports identified by the dedicated ranking approach, the retriever focuses on extracting content related to the medical query. We here apply existing text retrievers, such as BM25 and Contriever, during this phase.
@@ 3.3 EHR-Based Report Ranker $f_{\text{Ranker}}$
Efficiency Analysis The overall computation is practically efficient since the computation of Jaccard Index can be significantly accelerated with some libraries, such as Faiss Douze et al. (2024) and NumPy Harris et al. (2020). Besides, indexing medical entities from tabular data enables fast lookups, further reducing computational overhead.
Task Answer Format #Query Example Question Practical Significance Background Source Option Source
@@ 4.1 Dataset Introduction
Overview As shown in Table 1, DischargeQA consists of a total of 1,280 QA pairs, each associated with a patient ID and corresponding clinical background. The questions in DischargeQA can be categorized into three main types:
@@
Evaluation Settings Each problem in DischargeQA includes a problem description, the patient’s clinical background, and multiple options for the LLM to choose from. While instruction inference uses a single-select setup, diagnosis and medication inference adopt a multi-select setup, requiring the model to select multiple options to answer the questions. Each option corresponds to a specific diagnosis or treatment. This multi-select format presents a more challenging and realistic setting for LLMs, as clinicians often need to identify and address multiple coexisting conditions.
@@ 4.2 Dataset Construction
Patients Filtering We first filter out low-quality patient records in MIMIC-IV for various reasons. Starting with all 430,000 patients, we remove encounters without available discharge summaries, leaving 320,000 patients. Next, we filter out patients with fewer than 3 or more than 40 entries in any of the diagnosis, medication, or procedure records, resulting in a final dataset of 28,000 patients for generating QA pairs. For instruction inference, we further exclude patients with excessively short discharge summaries using GPT-4o.
Background Generation To enable LLMs to make realistic medical decision, it is necessary to offer a clinical background of the patient along with the question. To avoid label leakage during the context generation, we propose leveraging the structured format of the discharge summary, which consists of the following components:
@@
Option Generation For all tasks, the correct options are directly extracted from the discharge reports. For diagnosis and medication inference, we utilize EHR to generate non-trivial candidate options by extracting all associated diagnoses and medications of a patient and feeding them into GPT-4o to select contextually relevant candidates, ensuring a challenging selection process. For instruction inference, GPT-4o first summarizes the key points of the ground-truth answers and then applies permutations to generate plausible yet incorrect candidate options.
@@ 5 Experiments
In this section, we evaluate ExpRAG on DischargeQA with four LLMs, comparing it with the text-based report ranker. We also analyze the effects of balancing coefficients, the number of similar patients, and retriever choices, followed by case studies on the retrieved experience.
Model Report Ranker Instruction Diagnosis Medication
Acc(%) Acc(%) F1 Acc(%) F1
GPT-3.5 Direct-Ask 73.0 18.50 0.498 1.15 0.234
Text-based 78.8 15.60 0.405 0.68 0.317
EHR-based 79.5 18.81 0.504 1.80 0.371
GPT-4o Direct-Ask 90.0 9.86 0.510 3.65 0.486
Text-based 90.3 8.26 0.493 4.95 0.601
EHR-based 91.3 21.33 0.530 9.68 0.638
Deepseek-R1 Direct-Ask 70.0 10.78 0.363 0.69 0.217
Text-based 72.0 12.84 0.381 2.03 0.230
EHR-based 75.3 11.01 0.379 1.13 0.241
Mistral-7b Direct-Ask 67.0 16.03 0.488 1.13 0.362
Text-based 70.0 14.67 0.490 0.90 0.344
EHR-based 69.0 13.79 0.505 1.13 0.353
@@ 5.1 Comparison of LLMs
We evaluated the performance of 4 state-of-the-art LLMs of varying scales, ranging from close-source to 8 billion parameters open-source model, on three clinical tasks: discharge instructions, diagnosis, and medications. These models included (1) Proprietary LLMs GPT-3.5 OpenAI (2022) and GPT-4o OpenAI (2024)) (2) Open-source LLMs Mistral Jiang et al. (2023a), and Deepseek-R1 Guo et al. (2025).
Hyperparameters We also present the results of LLMs with ExpRAG. The default balancing coefficients $\lambda_{1/2/3}$ are set to 1/3 each and auto-merging Liu (2022) is used as the retriever. The number of similar patients is set as 15.
Metrics We report accuracy across all the tasks, calculated as the percentage of correctly answered questions. Note that for multi-select problems, a question is considered answered correctly only when all correct options are selected. We additionally report F1 scores for the multi-select problems to provide a more comprehensive analysis of LLM performance on these challenging tasks.
Results Table 2 summarizes the performance of all LLMs with ExpRAG on the three clinical tasks. GPT-4o consistently achieves the best performance across tasks, with an absolute improvement of 13% over GPT-3.5 on instruction inference problems. Additionally, the multi-select tasks (diagnosis and medication) prove significantly more challenging, as most LLMs achieve accuracy below 20%, indicating the limited medical reasoning capabilities of current large language models.
@@ 5.2 Comparison of Report Ranker
To verify the effectiveness of ExpRAG and the utilization of EHR, we introduce a baseline that performs report ranking solely based on text, i.e., a text-based ranker. We embed all text components, including the question, options, and background, into a single representation. Similarly, each discharge report is embedded, and similarity is computed with the question embedding. The top-$k$ reports are then retrieved, followed by a text retriever to extract the relevant information.

### v2

@@ Abstract
To improve the reliability of Large Language Models (LLMs) in clinical applications, retrieval-augmented generation (RAG) is extensively applied to provide factual medical knowledge. Beyond general medical knowledge, clinical case-based knowledge is also critical for effective medical reasoning, as it provides context grounded in real-world patient experiences. Motivated by this, we propose Experience Retrieval-Augmentation (ExpRAG) framework based on Electronic Health Record (EHR), aiming to offer the relevant context from other patients’ discharge reports. ExpRAG performs retrieval through a coarse-to-fine process: it first applies an EHR-based report ranker to efficiently identify similar patients as experience, and then utilizes a context retriever to extract task-relevant content for enhanced medical reasoning. To evaluate RAG systems on EHR data including ExpRAG and medical agents, we introduce DischargeQA, a clinical QA dataset with 1,280 discharge-related questions across diagnosis, medication, and instruction tasks. Each problem is generated using historical EHR data to ensure realistic and challenging scenarios. Experimental results demonstrate that ExpRAG consistently outperforms traditional text-based rankers, achieving an average relative improvement of 5.2%, highlighting the importance of case-based knowledge for medical reasoning.
@@ 1 Introduction
Figure 1: An illustrative example of utilizing experience from relevant clinical cases to support medical decision: adjusting a patient’s warfarin dosage based on the specific clinical context rather than relying on a generic standard dose
@@
In light of this, we argue that, in addition to general factual concepts, clinical case-based knowledge is also crucial for effective medical reasoning. The intuition is that an experienced clinician often relies on past cases with similar conditions to guide diagnosis, treatment decisions, and discharge planning. To this end, we propose Experience Retrieval-Augmentation (ExpRAG) framework, leveraging a large-scale EHR database MIMIC-IV Johnson et al. (2023b) as its knowledge basis. Specifically, ExpRAG breaks down the retrieval process into two steps: (1) report ranking applies an EHR-based similarity measurement to identify patients with similar medical conditions, and (2) experience retrieval extracts problem-relevant content from these patients’ discharge reports, which serves as the case-based contextual knowledge for LLMs. The introduced EHR modality enables large-scale clinical experience retrieval, grounding the model’s reasoning in real-world clinical practices.
To evaluate capability of ExpRAG and other RAG methods/agents in medical reasoning, we introduce DischargeQA, a clinical dataset including 1,280 QA pairs dedicated to discharge-related problems. The dataset primarily includes three types of problems:simulating the discharge process of final diagnosis, medication prescription and post-discharge instructions. For each problem, we follow the data structure of the discharge report and select the content preceding the question as the problem background to avoid label leakage. Additionally, we index the option candidates using EHR to generate contextually relevant options, ensuring a non-trivial and clinically meaningful challenge for the model.
We evaluate the performance of five different LLMs and compare ExpRAG with the text-based report ranker using DischargeQA.  [^1] Uploading to https://physionet.org/ , please check updates on Github: https://github.com/jou2024/EXPRAG
The results demonstrate the effectiveness of using EHR to retrieve relevant clinical experience, as it consistently improves the performance of LLM backbones and outperforms the text-based ranker with an average relative improvement of 5.2%. Our main contributions are summarized as follows:
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG has become a key paradigm for overcoming the static knowledge limitations of LLMs by retrieving external information Gu et al. (2018); Petroni et al. (2019). Traditional RAG frameworks typically use dense retrieval methods to augment generative tasks Devlin et al. (2019); Xiong et al. (2021). While effective in general QA, these methods often lack domain specificity, which is critical in healthcare Lu et al. (2024). Recent advancements like ClinicalRAG Lu et al. (2024) and MIRAGE Xiong et al. (2024) address this by integrating structured EHR data and clinical notes for diagnosis and treatment planning. However, existing benchmarks primarily focus on isolated information retrieval Johnson et al. (2023a); Kweon et al. (2024), overlooking the complexities of reasoning over patient histories and similar cases. To bridge this gap, our work extends RAG by combining structured EHR data with discharge summaries, enabling experience-driven reasoning for more realistic and reliable medical QA.
Medical QA Benchmark. EHRSQL (Lee et al., 2022) and DrugEHRQA (Bardhan et al., 2022) target structured data queries, with the former addressing SQL-based operations and the latter focusing on drug-related questions. EHRNoteQA (Kweon et al., 2024) and RadQA (Soni et al., 2022) leverage clinician-verified QA pairs from discharge summaries and radiology reports, while MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019) evaluate LLMs with questions from medical exams or PubMed articles. Discharge-summary-focused datasets like emrQA (Pampari et al., 2018) and CliniQG4QA (Yue et al., 2021) use discharge notes for QA tasks, and specialized datasets like RxWhyQA (Fan, 2019) and drug-reasoning QA (Moon et al., 2023) focus on specific question types like medication reasoning. Unlike these benchmarks, DischargeQA introduces an evaluation framework centered around the discharge process, simulating the clinical workflow from diagnosis inference to medication prescription and discharge instruction generation. Additionally, we leverage EHR to generate non-trivial, contextually relevant candidate options, providing a more challenging and realistic setup.
@@ 3 ExpRAG Framework
ExpRAG provides a comprehensive framework for retrieving relevant knowledge from the cohort, as shown in Figure 2. In this section, we first formulate the problem that ExpRAG aims to tackle and then elaborate the two-step retrieval framework.
@@ 3.1 Task Formulation
3.2 Coarse-to-Fine Retrieval Framework.
@@
Report Ranking. Before addressing a specific medical query, an intuitive assumption is that only patients with similar clinical histories, e.g., similar diseases or medications, can potentially provide meaningful guidance. In light of this, ExpRAG first employs a report ranker to efficiently discard unrelated cases and narrow down the candidate pool using the patient information:
@@
where $\mathcal{D}^{\prime}=\{D_{i}\}_{N^{\prime}\ll N}$ is a small subset of the selected discharge summaries. The ranker module will be scalable and enable effective utilization of patient context. Specifically, we introduce EHR as a knowledge base to facilitate the patient-level similarity measurement, as presented in Section 3.3.
Experience Retrieval. Based on the selected candidate pool, a sophisticated text retriever is capable of providing more accurate and dedicated clinical experience searching:
@@
Built on top of clinically relevant reports identified by the dedicated ranking approach, the retriever focuses on extracting content related to the medical query. We here apply existing text retrievers, such as auto-merging or BM25, during this phase.
@@ 3.3 EHR-Based Report Ranker $f_{\text{Ranker}}$
Efficiency Analysis. The overall computation is practically efficient since the computation of Jaccard Index can be significantly accelerated with some libraries, such as Faiss Douze et al. (2024) and NumPy Harris et al. (2020). Besides, indexing medical entities from tabular data enables fast lookups, further reducing computational overhead.
Task Response Type #Query Example Question Practical Significance Background Source Option Source
@@ 4.1 Dataset Introduction
Overview. As shown in Table 1, DischargeQA consists of a total of 1,280 QA pairs, each associated with a patient ID and corresponding clinical background. The questions in DischargeQA can be categorized into three main types:
@@
Evaluation Settings. Each problem in DischargeQA includes a problem description, the patient’s clinical background, and multiple options for the LLM to choose from. While instruction inference uses a single-select setup, diagnosis and medication inference adopt a multi-select setup, requiring the model to select multiple options to answer the questions. Each option corresponds to a specific diagnosis or treatment. This multi-select format presents a more challenging and realistic setting for LLMs, as clinicians often need to identify and address multiple coexisting conditions.
@@
Model Context Instruction Diagnosis Medication
Acc(%) Acc(%) F1 Acc(%) F1
Mistral-7b Direct-Ask 67.0 16.03 0.488 1.13 0.362
bge-small-en 70.0 14.67 0.490 0.90 0.356
ExpRAG ${}_{\text{EHR}}$ 69.0 13.79 0.505 1.13 0.371
Deepseek-R1-8B Direct-Ask 70.0 10.78 0.363 0.69 0.217
bge-small-en 72.0 12.84 0.381 1.58 0.230
ExpRAG ${}_{\text{EHR}}$ 75.3 11.01 0.379 0.9 0.241
Qwen3-30B-A3B Direct-Ask 90.8 15.60 0.415 1.13 0.280
bge-small-en 93.8 18.12 0.502 2.25 0.366
ExpRAG ${}_{\text{EHR}}$ 95.3 17.43 0.528 1.35 0.355
GPT-3.5 Direct-Ask 73.0 18.50 0.498 1.15 0.234
bge-small-en 78.8 15.60 0.405 0.68 0.317
ExpRAG ${}_{\text{EHR}}$ 79.5 18.81 0.504 1.80 0.371
GPT-4o Direct-Ask 90.0 9.86 0.510 3.65 0.486
bge-small-en 90.3 8.26 0.493 4.95 0.601
ExpRAG ${}_{\text{EHR}}$ 91.3 21.33 0.530 9.68 0.638
@@ 4.2 Dataset Construction
Patients Filtering. We first filter out low-quality patient records in MIMIC-IV for various reasons. Starting with all 430,000 patients, we remove encounters without available discharge summaries, leaving 320,000 patients. Next, we filter out patients with fewer than 3 or more than 40 entries in any of the diagnosis, medication, or procedure records, resulting in a final dataset of 28,000 patients for generating QA pairs. For instruction inference, we further exclude patients with excessively short discharge summaries using GPT-4o, as explained in A.1.
Background Generation. To enable LLMs to make realistic medical decision, it is necessary to offer a clinical background of the patient along with the question. To avoid label leakage during the context generation, we propose leveraging the structured format of the discharge summary, which consists of the following components:
@@
Option Generation. For diagnosis and medication inference, we utilize EHR to generate non-trivial candidate options by extracting all associated diagnoses and medications of a patient and feeding them into GPT-4o to select contextually relevant candidates, ensuring a challenging selection process. For instruction inference, GPT-4o first summarizes the key points of the ground-truth answers and then applies permutations to generate plausible yet incorrect candidate options. Notably, EHR tabular data usually has very limited overlap with our task for discharge, such as in-hospital medications, due to professionalism and use cases, have much limitation than discharge prescriptions. For all tasks, the correct options are directly extracted from the discharge reports.
@@ 5 Experiments
In this section, we evaluate ExpRAG on DischargeQA with five LLMs, comparing it with the text-based report rankers. We also analyze the effects of balancing coefficients, the number of similar patients, and retriever choices, followed by case studies on the retrieved experience.
@@ 5.1 Comparison of LLMs
We evaluated the performance of 4 state-of-the-art LLMs of varying scales, ranging from close-source to 8 billion parameters open-source model, on three clinical tasks: discharge instructions, diagnosis, and medications. These models included GPT-3.5 OpenAI (2022), GPT-4o OpenAI (2024)), Mistral-7B Jiang et al. (2023a), and two thinking models—Deepseek-R1-8B Guo et al. (2025) and Qwen3-30B-A3B Team (2025).
Hyperparameters. We also present the results of LLMs with ExpRAG. The default balancing coefficients $\lambda_{1/2/3}$ are set to 1/3 each and auto-merging Liu (2022) is used as the retriever. The number of similar patients is set as 15.
Metrics. We report accuracy across all the tasks, calculated as the percentage of correctly answered questions. Note that for multi-select problems, a question is considered answered correctly only when all correct options are selected. We additionally report F1 scores for the multi-select problems to provide a more comprehensive analysis of LLM performance on these challenging tasks.
Results. Table 2 summarizes the performance of all LLMs with ExpRAG on the three clinical tasks. GPT-4o consistently achieves the best performance across tasks, with an absolute improvement of 13% over GPT-3.5 on instruction inference problems. Additionally, the multi-select tasks (diagnosis and medication) prove significantly more challenging, as most LLMs achieve accuracy below 20%, indicating the limited medical reasoning capabilities of current large language models, the value of challenge of DischargeQA.
@@ 5.2 Comparison of Report Ranker
To verify the effectiveness of ExpRAG and the utilization of EHR, we implement baselines that perform report ranking solely based on text, i.e., a text-based ranker, as a key part in other traditional RAG methods. Using embedding model bge-small-en-v1.5 Xiao et al. (2023), a popular model due to its light weight and high correlation scores in Table 4, we embed queries (question, options, background) to be computed similarity with each discharge report embedding. The top-$k$ similar reports are then retrieved, followed by a text retriever to extract the relevant information.
@@
As shown in Table 3, we compare our EHR-based ranker with two additional sentence-embedding models: all-MiniLM-L6-v2 and paraphrase-MiniLM-L3-v2 Reimers and Gurevych (2019), using GPT3.5 as backbone LLM. As a result, ExpRAG ${}_{\text{EHR}}$ outperforms all text-based variants in terms of all the metrics.
Rankers Instruction Diagnosis Medication
Acc Acc F1 Acc F1
bge-small-en 78.8 15.60 0.405 0.68 0.317
all-MiniLM-L6 79.3 17.43 0.476 1.13 0.316
paraphrase-L3 77.5 18.81 0.492 1.58 0.330
ExpRAG ${}_{\text{EHR}}$ 79.5 18.81 0.504 1.80 0.371
Table 3: Performance Comparison with Embedding Models as Reranker vs EHR-based Reranker.
Rankers Pearson Spearman
bge-small-en-v1.5 0.639 0.623
all-MiniLM-L6 0.640 0.618
paraphrase-MiniLM-L3 0.478 0.481
ExpRAG ${}_{\text{EHR}}$ 0.669 0.648
Table 4: Retrieval Performance Comparison of Rerankers using Pearson and Spearman Correlation.
Retrieval Correlation Comparison. We conducted an experiment comparing patients similarity scores generated by the EHR-based method with those reliably annotated by LLMs. A higher correlation between the two indicates stronger retrieval performance by the EHR-based approach. Specifically, we randomly sample 100 target patients from DischargeQA, as detailed in Appendix A.7. Each target–candidate pair is scored by GPT4o-mini on the three modalities (diagnoses, procedures, prescriptions) with a single “overall” similarity. For each ranker (i.e., three text-embedding models and our EHR-based ExpRAG), we compute Pearson and Spearman correlation between its scores and annotations, and average over all 100 targets (Table 4). Over the strongest text-embedding baselines,the results by ExpRAG confirm that explicit EHR similarity provides more faithful retrieval signals than generic sentence embeddings.
Figure 4: Comparison of Similar Patients.